OpenStudy (anonymous):

Find the Mistake: Provided below is a solution to the equation 10x + 10 − 14x + 8 = −13 + 55. Choose the statement that is true about the given solution. Line # 10x + 10 − 14x + 8 = −13 + 55 1 −4x + 18 = 42 2 −4x = 24 3 x = 6

OpenStudy (anonymous):

Well, simplify the equation and find out. 10x+10-14x+8=-13+55 Start by combining like terms: -4x+18=42 then subtract 18: -4x=24 All the sudden you find the answer.

OpenStudy (anonymous):

I didn't quite finish this I guess, but if you simplify one more time, it equals: x=-6

OpenStudy (anonymous):

ok thanx

OpenStudy (anonymous):

don't forget to medal? =P

OpenStudy (anonymous):

i want